OpenStudy (anonymous):

the area of a rectangle is 360 ft square. the perimeter is 76 feet. find the dimensions of the rectangle.

OpenStudy (tylerd):

360 = L x W 2L + 2W = 76

OpenStudy (tylerd):

2W = 76 - 2L W = 38 - L

OpenStudy (tylerd):

360 = L (38 - L) -L^2+38L=360

OpenStudy (tylerd):